OpenStudy (anonymous):

-5a+7b=-1 3a+2b= 13 show work please

OpenStudy (campbell_st):

multiply the 1st equation by 3 and the 2nd equation by 5 and add the resulting equations to eliminate a -15a + 21b = -3 15a + 10b = 65 ----------------- 31b = 62 b = 2 substitute b = 2 into either equation to find a

OpenStudy (anonymous):

multiply equation 1 from 2 and equation 2 from 7.. so equation 1= -10a+14b=-2 equation 2=21a+14b=91 subtract equations u will get value of b -31a=-93 a=3 now substitute a in equation 1 original..i.e=-5a+7b=-1 -15+1=-7b b=2

OpenStudy (anonymous):

could you do a check step?

OpenStudy (anonymous):

 hence proved..:)